40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 1 00:13:21 2014, Koji, Update, LSC, ALS X beat setup aligned 
    Reply  Fri May 2 13:03:30 2014, Jenne, Update, LSC, ALS Y beat setup aligned 
       Reply  Fri May 2 14:31:26 2014, Koji, Update, LSC, ALS Y beat setup aligned 
          Reply  Fri May 2 19:03:13 2014, Jenne, Update, LSC, ALS out of loop spectra ALS_OutOfLoop2_2May2014.pdf
Message ID: 9906     Entry time: Fri May 2 19:03:13 2014     In reply to: 9905
Author: Jenne 
Type: Update 
Category: LSC 
Subject: ALS out of loop spectra 

I have taken out of loop spectra for both arms, by looking at POX/POY while the arms were controlled with ALS.

To do this, I put the POY11 Q signal directly to the whitening board, which meant that I removed the cable coming from the common mode board.  (Now that we're doing CM stuff again, I have put it back, so POY is still in the slightly weird "going through the CM slow path" situation). 

For the locking, both arms had FMs 1, 2, 3, 5, 6 engaged.  Yarm had a gain of +17, Xarm had a gain of -17. 

Y beatnote was 98.6MHz with a peak height of -22 dBm.  X beatnote was 45.0MHz with a peak height -11 dBm.

I drove ITMY at 503.1 Hz with 100 counts.  I drove ITMX at 521.1 Hz with 25 cnt. 

Koji helped me match up the peak heights between the FINE_PHASE_OUT_HZ calibrated signals and the PDH signals. 

The out of loop noise is definitely below 1kHz rms now, which is better than it was!  Hooray!

ALS_OutOfLoop2_2May2014.pdf

ELOG V3.1.3-